# NuwaTS: a Foundation Model Mending Every Incomplete Time Series

### v2

@@ 4.5 Enhancing Forecasting on Incomplete Time Series
Table 7: We trained and evaluated TimesNet [38] on incomplete time series with original missing rates of 0.2 and 0.5, comparing its forecasting performance on data imputed by PatchTST and NuwaTS. We set the length of input sequence and forecasting both to 96. A lower MSE or MAE indicates a better forecasting performance. Red: the best.
@@
Missing Rate: 0.2 Default 0.457 0.450 0.353 0.389 0.389 0.404 0.200 0.284
@@
Missing Rate: 0.5 Default 0.585 0.515 0.370 0.403 0.572 0.496 0.238 0.318
@@
In practical applications, forecasting models often have to train on incomplete data, which can significantly impair their performance. To address this issue, we applied pre-trained imputation models to impute the incomplete data before training the forecasting model. We chose TimesNet [38] as the forecasting model and used four ETT datasets for training. We randomly discarded 20% and 50% of the data in the training sets of four ETT datasets.
To prevent data leakage during imputation, we used NuwaTS and PatchTST pre-trained on LargeST to impute the training data. Finally, we trained TimesNet [38] on the imputed time series and tested it on the complete time series.
@@ 4.6 Fine-tuning Imputation Model to Forecasting Model
Table 8: Forecasting results. We set the length of input sequence and forecasting both to 96. Forecasting results from other baselines come from  [38; 20; 5]. We fine-tuned the model pre-trained on LargeST in order to avoid data leakage. A lower MSE or MAE indicates a better performance. Red: the best, Blue: the second best.
@@
NuwaTS(fine-tuned without output layer) 0.398 0.419 0.319 0.369 0.316 0.357 0.191 0.275 0.197 0.284 0.176 0.219
NuwaTS(fine-tuned with output layer) 0.375 0.404 0.292 0.348 0.319 0.360 0.185 0.268 0.183 0.267 0.172 0.215
By directly inserting the masked padding tokens $\mathbf{p}$ where $\mathbf{p}\in\mathbb{R}^{M\times D}$ after the original input embeddings $\mathbf{E}_{i}\in\mathbb{R}^{(N+2)\times D}$, we get $\mathbf{E}_{i}^{f}=[\mathbf{k},\mathbf{z}_{i,(v_{g})},\mathbf{E}_{i,(p)},\mathbf{p}]$ where $\mathbf{E}_{i}^{f}\in\mathbb{R}^{(N+2+M)\times D}$. The model automatically imputes the future $M$ patches, effectively transforming into a forecasting model. We discard the prefixal part and obtain the final forecasting represenation $\mathbf{h}^{(\text{Layer})}_{i}\in\mathbb{R}^{M\times D}$. Then $\mathbf{h}^{(\text{Layer})}_{i}$ pass through the output layer and we get the final forecasting results. In order to further help NuwaTS adapt to the forecasting task, we applied domain-specific fine-tuning on the one-for-all model and trained its domain-transfer layer, LayerNorm layers as well as position embeddings, which account for only 9.35% of the model’s total parameters (details in Appendix A.3). Notably, even without training the output layer, the model still achieves remarkable performance on the long-term forecasting task, as shown in Table 8. We visualized the forecasting results in Appendix C.2.
@@ 5 Conclusion
We propose NuwaTS, a time series imputation model equipped with a PLM backbone and a contrastive learning module that explicitly incorporates patch-wise statistical information and missing patterns. Experimental results demonstrate that our imputation model significantly outperforms existing methods and can adapt to diverse domains and missing rates in a zero-shot manner. Additionally, we developed an efficient domain-specific fine-tuning technique that, with minimal cost, can adapt the model from an all-in-one to a domain-specific one, and also transform it from an imputation model to a forecasting model. To the best of our knowledge, this is the first attempt to design a foundational time series imputation model that can be applied across different domains. Our proposed method serves as a valuable reference for this field.
@@ A.1 Dataset Details
Our research encompass experimental evaluations utilizing a selection of ten popular multivariate datasets, including (1) ETT [^5] https://github.com/zhouhaoyi/ETDataset reports seven factors of electricity transformers, encompassing four subsets (ETTm1, ETTm2, ETTh1, ETTh2). These are divided into two categories based on temporal granularity, with data recorded at intervals of 15 minutes (m) and 1 hour (h), respectively. (2) ECL [^6] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 records the hourly electricity consumption of 321 customers. (3) Weather [^7] https://www.bgc-jena.mpg.de/wetter/ comprises 21 meteorological factors, recorded at ten-minute intervals throughout the year 2020, including variables such as temperature, humidity. (4) PEMS encompasses four datasets (03, 04, 07 and 08), each of which collects data on California’s public traffic network at a frequency of every five minutes. For the LargeST [19], which consists of 8,600 sensors and spans from 2017 to 2021, we only selected the 2019 data for pre-training NuwaTS in this paper.
@@
Using the same approach, we sliced LargeST into 100.1 million segments to train NuwaTS and PatchTST. This was done for zero-shot evaluation and fine-tuning for forecasting tasks on ETT, weather and ECL, ensuring no data leakage.
@@ A.2 Implementation Details
Regarding the training details of our model compared to others, we mainly followed the imputation task configuration including optimizer, learning rate and early stop strategy in the https://github.com/thuml/Time-Series-Library for fair comparisons. We trained three different PLM backbones: GPT2 [30], BERT [7], and LLaMA [35]. All the comparison models, as well as GPT2 and BERT, were trained on NVIDIA RTX 3090-24G GPUs. LLaMA-version (the first six layers) was trained on an NVIDIA A6000-48G GPU.
@@ A.3 Different parameter-efficient training strategy
Table 9: Different parameter-efficient training strategy on GPT2, BERT and LLaMA2.
[b] Task InputSize Backbone(number of layers) TrainingModule MemoryOccupation(GB) Training/TotalParameter(M) Imputation (512,96,1) GPT2(3) LN, WPE, IN&OUT, FFN 2.7 20.3/66.0 Imputation (512,96,1) GPT2(6) LN, WPE, IN&OUT, FFN 4.6 38.0/90.8 Imputation (512,96,1) GPT2(9) LN, WPE, IN&OUT, FFN 6.4 55.8/115.6 Imputation (512,96,1) GPT2(12) LN, WPE, IN&OUT, FFN 8.3 73.5/140.4 Imputation (512,96,1) LLaMA2(6) LN, IN&OUT 19.5 6.3/1351.7 Imputation (128,96,1) LLaMA2(6) LN, IN&OUT, FFN 20.3 818.0/1351.7 Imputation (512,96,1) BERT(6) LN, WPE, IN&OUT, FFN 3.1 34.1/68.2 Imputation(fine-tuned) (512,96,1) GPT2(6) DTL 2.4 7.7/90.8 Forecasting(w/o-OutputLayer) (512,96,1) GPT2(6) DTL, LN, WPE 3.8 8.5/90.8 Forecasting(w-OutputLayer) (512,96,1) GPT2(6) DTL, LN, WPE, OUT 3.8 8.9/90.8
@@
  • FFN: Feed-Foward Network, DTL: Domain-Transfer Layer&prefix
@@ B.1 Algorithm
Algorithm 2 Domain Specific Fine-tuning
1:  Initialize continuous prefix $\boldsymbol{\mathcal{P}}\in\mathbb{R}^{2\times\text{Layer}\times d}$
2:  Obtain $\boldsymbol{\mathcal{\hat{K}}}\in\mathbb{R}^{2\times\text{Layer}\times d}\leftarrow\text{DomainTrans}\left(\mathbf{k}\right)$
@@
6:     Obtain $[\mathbf{Key}_{p},\mathbf{Value}_{p}]\leftarrow\boldsymbol{\mathcal{P}}^{(n)}+\beta\boldsymbol{\mathcal{\hat{K}}}^{(n)}$ where $\boldsymbol{\mathcal{P}}^{(n)}\in\mathbb{R}^{2\times d}$
@@ B.2 Instance Norm
We applied reversible instance normalization (RevIN) [14] to individual series before splitting them into patches. Each series was normalized to have zero mean and unit standard deviation, with the original mean and standard deviation stored for later reversal, thereby reducing the domain distribution shift caused by non-stationary information. The details are as follows:
@@ B.3 Contrastive Learning Loss
Given a time patch represenation as a query $q$, and a set of keys $\mathbb{K}=\left\{k_{0},k_{1},\ldots\right\}$, including a positive target $k_{+}$. We employ the InfoNCE [27] as a auxiliary loss function to optimize the model’s performance:
@@
where we employ bi-linear inner-product and $W$ is a learnable matrix.
@@ C.2 Forecasting Visualization
We visualized the forecasting results of fine-tuned NuwaTS whose backbone had not undergone any forecasting training in Figure 10.
((a)) ECL
((b)) ECL
((c)) ECL
((d)) ETTh1
((e)) ETTh1
((f)) ETTh1
((g)) ETTh2
((h)) ETTm1
((i)) ETTm2
((j)) Weather
Figure 10: Case visualization when sequence length set to 96 and forecasting length set to 96.
@@ C.3 Domain Recognition
We also visualized the domain recognition ability of NuwaTS in Figure 11. By directly appending domain-specific embeddings $\mathbf{k}$ again to the end of the input embedding $\mathbf{E}_{i}\in\mathbb{R}^{(N+2)\times D}$, we get $\hat{\mathbf{E}_{i}}=[\mathbf{k},\mathbf{z}_{i,(v_{g})},\mathbf{E}_{i,(p)},\mathbf{k}]$, where $\hat{\mathbf{E}_{i}}\in\mathbb{R}^{D\times(N+3)}$. We feed the $\hat{\mathbf{E}_{i}}$ into the NuwaTS (one-for-all) model, and get the final representation $\hat{\mathbf{h}}^{(\text{Layer})}_{i}\in\mathbb{R}^{(N+3)\times D}$. Then We extract the last embeddings $\mathbf{k^{\prime}}\in\mathbb{R}^{D}$ from $\hat{\mathbf{h}}^{(\text{Layer})}_{i}$ which is mixed with the patch embeddings in front of the input embeddings, thus exhibiting distinct domain characteristics. We collected $N$ time seires from different domain and extracted $\mathbf{k^{\prime}}$. After applying the t-SNE [37] method to reduce the dimension to 2, we obtained scattered points $\mathbf{k^{\prime}}^{p}\in\mathbb{R}^{N\times 2}$ and visualized them. The scattered points corresponding to time series from the same domain tend to cluster together which indicates that NuwaTS could recognize domain information.
@@ Appendix D Extra Experiments
D.1 Learning From Incomplete Training data

### v3

@@ 4.6 Enhancing Forecasting on Incomplete Time Series
Table 7: We trained and evaluated TimesNet (Wu et al., 2022) on incomplete time series with original missing rates of 0.2 and 0.5, comparing its forecasting performance on data imputed by PatchTST and NuwaTS. We set the length of input sequence and forecasting both to 96. A lower MSE or MAE indicates better forecasting performance.
@@
Missing Rate: 0.2
Default 0.457 0.450 0.353 0.389 0.389 0.404 0.200 0.284
@@
Missing Rate: 0.5
Default 0.585 0.515 0.370 0.403 0.572 0.496 0.238 0.318
@@
In practical applications, forecasting models often have to train on incomplete data, which can significantly impair their performance. To address this issue, we applied pre-trained imputation models to impute the incomplete data before training the forecasting model. We chose TimesNet (Wu et al., 2022) as the forecasting model and used four ETT datasets for training. We randomly discarded 20% and 50% of the data in the training sets of four ETT datasets.
To prevent data leakage during imputation, we used cross-domain NuwaTS and PatchTST pre-trained on LargeST to impute the training data. Finally, we trained TimesNet (Wu et al., 2022) on the imputed time series and tested it on the complete time series.
@@ 4.7 Fine-tuning Imputation Model to Forecasting Model
Table 8: Forecasting results. We set the length of input sequence and forecasting both to 96. Forecasting results from other baselines come from  (Wu et al., 2022; Liu et al., 2024c; Cai et al., 2024). We fine-tuned the cross-domain model in order to avoid data leakage. A lower MSE or MAE indicates a better performance.  Green : the best,  Yellow : the second best.
@@
NuwaTS(without inter-variable correlation)  0.375 0.404  0.292  0.348 0.319  0.360 0.185 0.268 0.183 0.267  0.172 0.215
NuwaTS(with inter-variable correlation)  0.374  0.403 0.308 0.357  0.314  0.357 0.184 0.267  0.151  0.242 0.179 0.221
By directly inserting the masked padding tokens $\mathbf{p}$ where $\mathbf{p}\in\mathbb{R}^{M\times D}$ after the original input embeddings $\mathbf{E}_{i}\in\mathbb{R}^{(N+2)\times D}$, we get $\mathbf{E}_{i}^{f}=[\mathbf{k},\mathbf{z}_{i,(v_{g})},\mathbf{E}_{i,(p)},\mathbf{p}]$ where $\mathbf{E}_{i}^{f}\in\mathbb{R}^{(N+2+M)\times D}$. The model automatically imputes the future $M$ patches, effectively transforming into a forecasting model. We discard the prefixal part and obtain the final forecasting represenation $\mathbf{h}^{(\text{Layer})}_{i}\in\mathbb{R}^{M\times D}$. Then $\mathbf{h}^{(\text{Layer})}_{i}$ pass through the output layer and we get the final forecasting results. In order to further help NuwaTS adapt to the forecasting task, we conducted two types of fine-tuning module based on the cross-domain model, the first one using two-layer MLP as the domain-transfer layer which accounts for only 9.35% of the model’s total parameters (details in Appendix A.3), and the other one computing the inter-variable correlation. (details in Appendix B.4) We visualized the forecasting results in Appendix C.2.
@@ 5 Conclusion
In this paper, we address the challenging problem of cross-variable and cross-domain generalization in time series imputation. We introduce NuwaTS, a time series imputation model built on a PLM backbone that explicitly captures patch-wise statistical information and missing patterns. Additionally, we propose an efficient domain-specific fine-tuning technique that allows the model to seamlessly adapt from an all-in-one imputation model to a domain-specific one with minimal cost, and even transform into a forecasting model when needed. To benchmark the model’s performance, we design a novel variable-wise train/validation/test partitioning strategy, providing a rigorous evaluation framework. Experimental results show that our approach significantly outperforms state-of-the-art methods, demonstrating strong adaptability across diverse domains and missing rates, even in zero-shot scenarios. To the best of our knowledge, this is the first foundational time series imputation model capable of generalizing across domains. We believe our method sets a new benchmark in the field and provides valuable insights for future research.
@@ A.1 Dataset Details
Our research encompasses experimental evaluations utilizing a selection of ten popular multivariate datasets, including (1) ETT [^3] https://github.com/zhouhaoyi/ETDataset reports seven factors of electricity transformers, encompassing four subsets (ETTm1, ETTm2, ETTh1, ETTh2). These are divided into two categories based on temporal granularity, with data recorded at intervals of 15 minutes (m) and 1 hour (h), respectively. (2) ECL [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 records the hourly electricity consumption of 321 customers. (3) Weather [^5] https://www.bgc-jena.mpg.de/wetter/ comprises 21 meteorological factors, recorded at ten-minute intervals throughout the year 2020, including variables such as temperature and humidity. (4) PEMS encompasses four datasets (03, 04, 07, and 08), each of which collects data on California’s public traffic network at a frequency of every five minutes. For the LargeST (Liu et al., 2024b), which consists of 8,600 sensors and spans from 2017 to 2021, we only selected the 2019 data for pre-training NuwaTS in this paper.
@@
Using the same approach, we sliced LargeST into 100.1 million segments to train NuwaTS and PatchTST. This was done for zero-shot evaluation and fine-tuning for forecasting tasks on ETT, weather, and ECL, ensuring no data leakage.
@@ A.2 Implementation Details
Regarding the training details of our model compared to others, we mainly followed the imputation task configuration, including optimizer, learning rate, and early stop strategy in the https://github.com/thuml/Time-Series-Library for fair comparisons. We trained three different PLM backbones: GPT2 (Radford et al., 2019), BERT (Devlin et al., 2019), and LLaMA (Touvron et al., 2023). All the comparison models, as well as GPT2 and BERT, were trained on NVIDIA RTX 3090-24G GPUs. LLaMA-version (the first six layers) was trained on an NVIDIA A6000-48G GPU.
@@ A.3 Different parameter-efficient training strategy
Table 9: Different parameter-efficient training strategies on GPT2, BERT, and LLaMA2.
[b] Task Input Size Backbone (number of layers) Training Module Memory Occupation (GB) Training / Total Parameter (M) Imputation (512,96,1) GPT2 (3) LN, WPE, IN&OUT, FFN 2.7 20.3 / 66.0 Imputation (512,96,1) GPT2 (6) LN, WPE, IN&OUT, FFN 4.6 38.0 / 90.8 Imputation (512,96,1) GPT2 (9) LN, WPE, IN&OUT, FFN 6.4 55.8 / 115.6 Imputation (512,96,1) GPT2 (12) LN, WPE, IN&OUT, FFN 8.3 73.5 / 140.4 Imputation (512,96,1) LLaMA2 (6) LN, IN&OUT 19.5 6.3 / 1351.7 Imputation (128,96,1) LLaMA2 (6) LN, IN&OUT, FFN 20.3 818.0 / 1351.7 Imputation (512,96,1) BERT (6) LN, WPE, IN&OUT, FFN 3.1 34.1 / 68.2 Imputation (fine-tuned) (512,96,1) GPT2 (6) DTL 2.4 7.7 / 90.8 Forecasting (w/o-OutputLayer) (512,96,1) GPT2 (6) DTL, LN, WPE 3.8 8.5 / 90.8 Forecasting (512,96,1) GPT2 (6) DTL, LN, WPE, OUT 3.8 8.9 / 90.8
@@
  • FFN: Feed-Forward Network, DTL: Domain-Transfer Layer&prefix
@@ B.1 Algorithm
Algorithm 1 One-for-all Model Training
1:  Initialize continuous prefix $\bm{\mathcal{P}}\in\mathbb{R}^{2\times\text{Layer}\times d}$
2:  Obtain $\bm{\mathcal{\hat{K}}}\in\mathbb{R}^{2\times\text{Layer}\times d}\leftarrow\text{DomainTrans}\left(\mathbf{k}\right)$
@@
6:     Obtain $[\mathbf{Key}_{p},\mathbf{Value}_{p}]\leftarrow\bm{\mathcal{P}}^{(n)}+\beta\bm{\mathcal{\hat{K}}}^{(n)}$ where $\bm{\mathcal{P}}^{(n)}\in\mathbb{R}^{2\times d}$
@@
Algorithm 2 Domain Specific Fine-tuning
1:  Initialize $\mathbf{X}\in\mathbb{R}^{N\times T}$, where $N$ is the number of variables and $T$ is the time steps.
2:  for $n=1$ to Layer do
3:     Map each variable $\mathbf{x}_{i}\in\mathbf{X}$ to a token: $\mathbf{z}_{i}^{(n)}\in\mathbb{R}^{d_{\text{light}}}$, resulting in a token sequence $\mathbf{Z}^{(n)}\in\mathbb{R}^{N\times d_{\text{light}}}$ for layer $n$
4:     Apply lightweight transformer at layer $n$ to capture inter-variable correlations:
5:     $\mathbf{h}^{(n)}\leftarrow\text{TransformerLayer}^{(n)}(\mathbf{Z}^{(n)})$, where $\mathbf{h}^{(n)}\in\mathbb{R}^{N\times d_{\text{light}}}$
6:     Pass the hidden state $\mathbf{h}^{(n)}$ through a linear layer:
7:     $\bm{\mathcal{\hat{K}}}^{(n)},\bm{\mathcal{\hat{V}}}^{(n)}\leftarrow\text{Linear}(\mathbf{h}^{(n)})$, where $\bm{\mathcal{\hat{K}}}^{(n)},\bm{\mathcal{\hat{V}}}^{(n)}\in\mathbb{R}^{N\times d_{\text{PLM}}}$
8:     $\bm{\mathcal{P}}^{(n)}\leftarrow\text{Concat}(\bm{\mathcal{\hat{K}}}^{(n)},\bm{\mathcal{\hat{V}}}^{(n)})$, where $\bm{\mathcal{P}}^{(n)}\in\mathbb{R}^{N\times 2\times d_{\text{PLM}}}$
9:  end for
10:  The generated prefix for all layers $\bm{\mathcal{P}}\in\mathbb{R}^{\text{Layer}\times N\times 2\times d_{\text{PLM}}}$ is then used for PLM fine-tuning.
Algorithm 3 Fine-tuning with inter-variable correlation
@@ B.2 Instance Norm
We applied reversible instance normalization (RevIN) (Kim et al., 2021) to individual series before splitting them into patches. Each series was normalized to have zero mean and unit standard deviation, with the original mean and standard deviation stored for later reversal, thereby reducing the domain distribution shift caused by non-stationary information. The details are as follows:
@@ B.3 Contrastive Learning Loss
Given a time patch represenation as a query $q$, and a set of keys $\mathbb{K}=\left\{k_{0},k_{1},\ldots\right\}$, including a positive target $k_{+}$. We employ the InfoNCE (Oord et al., 2018) as an auxiliary loss function to optimize the model’s performance:
@@
where we employ bi-linear inner product and $W$ is a learnable matrix.
B.4 Fine-tuning with inter-variable correlation
Unlike the two-layer MLP used in the channel-independent paradigm in Section 3.3, inspired by TimeXer (Wang et al., 2024b) and iTransformer (Liu et al., 2024c), we map each time series variable into a token. We then apply a stack of lightweight transformer layers, with the same number of layers as the PLM, to obtain the hidden states from each layer. These hidden states are passed through a linear layer to be directly mapped to $\bm{\mathcal{\hat{K}}}\in\mathbb{R}^{N\times 2\times\text{Layer}\times D}$, thus generating the prefix. This prefix now contains inter-variable correlation information, addressing the limitation of NuwaTS in modeling relationships between variables.
@@ C.2 Forecasting Visualization
We visualized the forecasting results of fine-tuned NuwaTS whose backbone had not undergone any forecasting training in Figure LABEL:forecasting_vis.
@@ C.3 Domain Recognition
We also visualized the domain recognition ability of NuwaTS in Figure 20. By directly appending domain-specific embeddings $\mathbf{k}$ again to the end of the input embedding $\mathbf{E}_{i}\in\mathbb{R}^{(N+2)\times D}$, we get $\hat{\mathbf{E}_{i}}=[\mathbf{k},\mathbf{z}_{i,(v_{g})},\mathbf{E}_{i,(p)},\mathbf{k}]$, where $\hat{\mathbf{E}_{i}}\in\mathbb{R}^{D\times(N+3)}$. We feed the $\hat{\mathbf{E}_{i}}$ into the NuwaTS (one-for-all) model, and get the final representation $\hat{\mathbf{h}}^{(\text{Layer})}_{i}\in\mathbb{R}^{(N+3)\times D}$. Then We extract the last embeddings $\mathbf{k^{\prime}}\in\mathbb{R}^{D}$ from $\hat{\mathbf{h}}^{(\text{Layer})}_{i}$ which is mixed with the patch embeddings in front of the input embeddings, thus exhibiting distinct domain characteristics. We collected $N$ time series from different domains and extracted $\mathbf{k^{\prime}}$. After applying the t-SNE (Van der Maaten & Hinton, 2008) method to reduce the dimension to 2, we obtained scattered points $\mathbf{k^{\prime}}^{p}\in\mathbb{R}^{N\times 2}$ and visualized them. The scattered points corresponding to time series from the same domain tend to cluster together, which indicates that NuwaTS can recognize domain information.
@@ Appendix D Extra Experiments
D.1 real-world dataset evaluation
Table 10: Imputation results on the Air-Quality dataset with missing values. 10% of the observed values in the validation set and test set are held out as ground truth for evaluation. A lower RMSE or MAE indicates better imputation performance.  Green : the best.
Model Air-Quality
RMSE MAE
GP-VAE 0.614 0.268
BRITS 0.525 0.153
SAITS 0.518  0.137
NuwaTS (zero-shot) 0.370 0.151
NuwaTS (domain fine-tuned)  0.363 0.144
We conducted experiments on Beijing Multi-Site Air-Quality (Zhang et al., 2017) in Table 10 following pipeline in SAITS (Du et al., 2023) where we observed that, despite not using inter-series correlations and the model not being directly trained on the data, the performance of the pre-trained one-for-all NuwaTS was close to that of SAITS, surpassing both BRITS and GP-VAE (Fortuin et al., 2020).
D.2 Ablation Study of Domain-Specific Fine-tuning
We also verified the design of domain-specific fine-tuning strategy of NuwaTS in Table 11. When the randomly initialized prefix key&value and the knowledge-transfer layer were ablated, the model’s fine-tuning performance showed a significant decline which indicates that domain-transfer layer retains the learned knowledge while newly added prefix key&value provide flexibility in transferring to specific domain. Additionally, we find that merely appending the prefix to the input layer and fine-tuning it significantly reduces the effectiveness of fine-tuning. Inserting the prefix into each layer of the PLM enhances the prefix’s representation capacity during domain transfer (Liu et al., 2022).
Table 11: Domain-specific fine-tuning ablation study on ETTs based on the model pre-trained on general fused dataset. A lower MSE or MAE indicates better imputation performance.  Green : the best.
Model ETTh1 ETTh2 ETTm1 ETTm2
MSE MAE MSE MAE MSE MAE MSE MAE
no fine-tuning 0.164 0.263 0.018 0.084 0.064 0.147 0.010 0.060
D.3 Learning From Incomplete Training data